# Towards AI-Safety-by-Design: A Taxonomy of Runtime Guardrails in Foundation Model based Systems

### v1

@@ IV-A2 Privacy
In FM-based systems, privacy is a critical concern due to the risks of handling large amounts of personal and sensitive data. Without proper guardrails, these systems can inadvertently expose or misuse this data [12, 1]. One key privacy risk is data leakage, where sensitive information is inadvertently revealed through model outputs. For instance, an FM-based LLM trained on private conversations might accidentally generate text that includes personal information such as names, addresses, or financial details [49]. This leakage can occur through direct responses or statistical inferences. For example, an FM-based customer service chatbot might inadvertently include sensitive information. In April-May 2023, a notable incident involved Samsung employees leaking proprietary information into ChatGPT, leading to Samsung banning ChatGPT [50].
@@ IV-A3 Security
Security in FM-based systems involves protecting them from malicious activities that could compromise their integrity and functionality [19, 14, 6]. For example, an FM-based system could be targeted by hackers to manipulate data, producing incorrect or harmful outputs that affect decision-making processes [51]. An incident similar to this reported in [52], where malicious users manipulated Microsoft’s Tay chatbot to produce inappropriate and offensive content, resulting in its shutdown. Additionally, FM-based systems are vulnerable to various threats [53]. Hackers can exploit vulnerabilities to access sensitive data, leading to breaches of confidentiality. Even with authorized access, there is a risk of data misuse by third-party providers [54]. FM-based systems are also prone to adversarial attacks, where specially designed queries extract sensitive information.
@@ IV-A4 Safety
FM-based systems face significant safety issues, particularly in generating harmful or misleading outputs. These issues can arise when models produce content that is inappropriate, offensive, or incorrect [3, 1]. For example, in applications where FM-based systems handle critical data, such as in medical diagnosis or or self-driving cars, incorrect outputs can lead to dangerous consequences [32]. Additionally, there is a risk of generating ethically or morally questionable content, which can damage the credibility and acceptance of the system [53].
@@ IV-A5 Fairness
FM-based systems can face bias and discrimination in model outputs. These biases can emerge from the training data, model algorithms, or deployment context [61, 2]. For instance, a language model used in recruitment for initial screening of CVs and responses to selection criteria might inadvertently favor candidates from certain demographics, cultures, and languages [26, 8], which ultimately undermines the credibility of the systems. Guardrails address these fairness issues and ensure equitable outcomes through model parameter adjustments. For instance, in our recruitment example, they can analyze the model’s decisions and correct potential biases. Furthermore, guardrails promote transparency and accountability by providing insights into how decisions are made [18, 1], and avoid harmful stereotypes [53].
@@ IV-A6 IP Protection
In FM-based systems, another key problem is the unauthorized use of generated content, where users might exploit outputs created by these systems without proper attribution or licensing, making them vulnerable to unauthorized use, duplication, or distribution [26, 55, 10]. Guardrails help mitigate these problems and better protect content’s copyright [49]. They also aid in compliance with copyright laws and licensing agreements to mitigate legal risks [19, 12]. Techniques such as watermarking, fingerprinting, and labeling ensure the copyright and ownership of data [10, 1]..
IV-A7 Compliance
Compliance in FM-based systems involves adhering to legal and regulatory standards [16, 20]. These issues are critical because non-compliance can lead to legal penalties, reputational damage, and loss of user trust. Guardrails reduce these risks by ensuring alignment with data protection regulations, industry standards, and guidelines [16, 54, 26]. They also facilitate better auditing and monitoring of the systems [18]. For instance, an FM-based system in a financial organization might use guardrails to monitor transactions, flag deviations from protocols, and ensure corrective actions are taken promptly. Additionally, guardrails assist in automating compliance checks and improving monitoring accuracy. They ensure that all aspects of the FM-based system’s operations align with the necessary legal and regulatory frameworks [62, 36], and better support internal audits and external regulatory reviews [12].
IV-B Quality Attributes Considered in the Design of Guardrails (RQ2)
We identified eight key quality attributes essential for the design of runtime guardrails in FM-based systems. Below, we discuss these attributes in detail, highlighting their roles and importance.
@@ IV-B3 Customizability
Customizable guardrails provide tailored protection that meets specific requirements and supports diverse operational needs in FM-based systems [1, 65]. They allow for adjustments and configurations that align with particular operational goals, data characteristics, and regulatory environments. For example, a customer service chatbot can enable priorities for different guardrails and adjust data handling based on the user’s location, ensuring compliance with GDPR in Europe and CCPA in the U.S. [66]. Customizability also includes the capability to integrate user-defined rules and policies, ensuring that the guardrails can support diverse operational contexts and evolving needs [64].
@@ IV-B6 Portability
Portability in guardrails for FM-based systems refers to the ability of these protective measures to be easily adapted and applied across different foundation model (FM) systems and platforms [27]. This includes ensuring that the guardrails function consistently across various FM architectures and environments, thereby maintaining their effectiveness and integrity regardless of the underlying system [26]. For example, the same guardrail can be applied for content moderation in both a customer service chatbot and a social media platform, regardless of their underlying technology. The benefits of designing portable guardrails include compatibility across multiple programming languages and frameworks facilitate their integration into diverse technological stacks [49]. These capabilities ensure that the guardrails remain effective and operational as the system evolves or migrates to new environments. Portable guardrails also support seamless updates and improve scalability to maintain high standards of security and compliance while adapting to new technological advancements within systems [16].
@@ IV-B8 Interpretability
IV-C Design of Guardrails (RQ3)
This section presents a structured taxonomy for designing guardrails, focusing on identifying various design alternatives.
@@ IV-C1 Guardrail Actions
Guardrail actions are crucial for addressing the specific needs of FM-based system components. We have identified the following guardrail actions as key elements for FM-based systems:
  • Block: The block action prevents specific inputs (such as user prompts) or outputs (such as content generated by FMs) from being processed or sent by various components (such as FMs and tools) in FM-based systems [54]. For example, the block action can reject the user prompts containing harmful instructions, thus preventing undesired outcomes.
  • Filter: The filter action involves scanning and removing undesired or irrelevant content from the inputs or outputs of different components in FM-based systems [69, 70]. For instance, a filter may remove any personal data contained in the user prompts or the output generated by FMs.
  • Flag: The flag action is used to mark specific inputs, outputs, operations within FM-based systems [16]. For example, unusual transactions requested by the FM-based agent can be flagged for human review to ensure they comply with organizational policies [1, 30].
  • Modify: The modify action allows for the adjustment of inputs or outputs of various components in the FM-based systems to meet specific requirements or standards [9]. For example, the system can modify the user prompts by adding more context and examples, making it easier for the FM to accurately interpret the user’s intentions and provide more relevant responses.
  • Validate: The validate action checks inputs, outputs, intermediate results against predefined criteria to ensure they meet specified requirements or standards [70, 26]. For example, the customer service chatbot could validate if the output generated by the FM-based system contains any sensitive company information.
@@
  • Simulate: The simulate action involves running tests in a controlled environment to predict and analyze potential outcomes [1]. For instance, a system might simulate different configurations to determine the most effective plan.
@@
FM-based Systems Prompts Block, Filter, Flag, Modify, Validate, Prioritize, Rate Limit, Parallel Calls, Retry, Defer, Isolate, Log
@@ IV-C3 Guardrails Scopes
The scope of guardrails in FM-based systems ranges from individual preferences to industry standards. At the user level, guardrails reflect individual preferences and requirements. This involves adjusting the system’s behavior based on user-defined settings to align outputs with both user expectations and ethical considerations. Incorporating user preferences into guardrails provides a personalized experience while maintaining safety and compliance [55, 69]. Such guardrails ensure that the system respects user autonomy and produces outputs that are relevant and acceptable.
At the organizational level, guardrails align with internal policies and procedures governing the operation and use of FM-based systems. This includes compliance with corporate governance, data protection policies, and ethical guidelines established by the organization [12]. Guardrails also ensure consistency and accountability across different departments and functions within the organization.
Industry-level regulations and standards provide the broader regulatory framework within which FM-based systems must operate. Guardrails designed to comply with these regulations guarantee that the system adheres to industry best practices and legal requirements [16]. Guardrails facilitate simpler auditing and certification processes, ensuring the system remains compliant with evolving regulatory landscapes.
@@ IV-C4 Rules
Guardrails rules can be configured in different ways: including uniform rules, priority-enabled rules, context-dependent rules, and negotiable rules. A uniform strategy applies the same set of guardrails consistently across all scenarios, ensuring simplicity and uniformity [67]. It is particularly effective in environments with stable and well-understood risks. It largely reduces the complexity of managing diverse guardrails [55]. A priority-enabled strategy prioritizes certain guardrails based on the criticality and sensitivity of operations or data. Context-dependent strategies adjust the implementation of guardrails based on the system’s specific operational context. This allows for dynamic adjustments to guardrails in response to changing conditions, user needs, and operational environments [49]. The negotiability of guardrails, categorized into hard and soft, defines the level of flexibility in enforcing rules. Soft guardrails allow adjustments based on context and situational demands, providing a balance between protection and operational flexibility [49]. In contrast, hard guardrails are rigid and non-negotiable, ensuring adherence to critical legal, ethical, or safety standards [12, 32].
@@ IV-C6 Modality
The modality of guardrails refers to the types of data and interactions they manage. Guardrails can be designed for single modal or multimodal systems. Single modal systems operate with one type of data input or output, such as text, image, or audio. For instance, in text-based systems, guardrails focus on addressing issues like offensive language, misinformation, and data privacy [49]. In image-based systems, they may involve techniques for detecting explicit content or ensuring image quality standards [26].
@@ IV-C7 Underlying Technique
Rule-based models utilize predefined rules to monitor and control FM-based system behavior. These models implement strict and deterministic guidelines that the system must follow to ensure compliance with regulatory requirements for data access and processing [49]. They are particularly effective in environments where operational parameters are well-defined and stable. Rule-based systems can be updated and are somewhat flexible. However, they may still struggle with unexpected scenarios, such as detecting novel AI-generated content that falls outside predefined rules. This reliance on static rules can limit their adaptability and regular updates are needed [71, 16].
@@
Hybrid models integrate rule-based systems with the adaptability of machine learning models to respond to new threats and evolving data patterns [53]. For instance, Khorramrouz et al.[59] demonstrate the use of the PaLM 2 framework to process user input and dynamically implement rule-based decisions. This framework tests the system’s limits by iteratively generating toxic content to evaluate PaLM 2’s safety guardrails. However, integrating hybrid models can increase system complexity and create additional challenges [53].
V Threats to Validity
@@
Our study is subject to standard literature search and selection bias threats. We addressed these threats by searching the most commonly used databases in the IT and software engineering domains. We adjusted our search strings several times during the automatic search to maximize the number of relevant articles matching two key concepts: ‘guardrails’ and ‘FM-Based systems’. We also kept our search string generic to search through the titles, abstracts, keywords, and full text of articles to cover the maximum number of relevant papers. We then conducted a manual search on Google Scholar to complement the automatic search using a snowballing strategy. Furthermore, predefined review protocols with detailed inclusion and exclusion criteria helped us reduce bias in selecting primary studies. We applied several quality assessment criteria (shown in section III-F) to estimate the quality of the selected primary studies. Even though the proposed criteria were not too strict, applying them indeed caused several initially selected papers to be excluded. To mitigate the risk of missing important data from the primary studies, we reinstated the excluded papers closely related to the primary studies.
@@
Moreover, our definitions and categorizations may not capture all relevant aspects of guardrails in FM-based systems. To mitigate this threat, we validated the taxonomy through extensive literature review and expert feedback. However, this introduces a risk of producing biased results that address only expert needs, as the people involved in the feedback process have extensive experience in the AI and software engineering domains. Our review protocols helped us to reduce such bias.
@@
We prepared a guardrails taxonomy and conducted a comparative analysis of its components to help the reader better understand their design and evaluation. We critically examined the strength and consistency of relationships in the selected studies for a reliable taxonomy structure and drew conclusions. Nonetheless, the generalizability of guardrails to different contexts and types in FM-based systems remains a potential limitation, and specific adaptations might be necessary for certain systems, such as those used in healthcare or financial organizations. Additionally, our findings face a challenge as some of them may become outdated due to the rapid evolution of FM-based systems and associated guardrails.

### v2

@@ IV-A2 Privacy
In FM-based agents, privacy is a critical concern due to the risks of handling large amounts of personal and sensitive data. Without proper guardrails, these systems can inadvertently expose or misuse this data [12, 1]. One key privacy risk is data leakage, where sensitive information is inadvertently revealed through model outputs. For instance, an FM-based LLM trained on private conversations might accidentally generate text that includes personal information such as names, addresses, or financial details [49]. This leakage can occur through direct responses or statistical inferences. For example, an FM-based customer service chatbot might inadvertently include sensitive information. In April-May 2023, a notable incident involved Samsung employees leaking proprietary information into ChatGPT, leading to Samsung banning ChatGPT [50].
@@ IV-A3 Security
Security in FM-based agents involves protecting them from malicious activities that could compromise their integrity and functionality [19, 14, 6]. For example, an FM-based system could be targeted by hackers to manipulate data, producing incorrect or harmful outputs that affect decision-making processes [51]. An incident similar to this reported in [52], where malicious users manipulated Microsoft’s Tay chatbot to produce inappropriate and offensive content, resulting in its shutdown. Additionally, FM-based agents are vulnerable to various threats [53]. Hackers can exploit vulnerabilities to access sensitive data, leading to breaches of confidentiality. Even with authorized access, there is a risk of data misuse by third-party providers [54]. FM-based agents are also prone to adversarial attacks, where specially designed queries extract sensitive information.
@@ IV-A4 Safety
FM-based agents face significant safety issues, particularly in generating harmful or misleading outputs. These issues can arise when models produce content that is inappropriate, offensive, or incorrect [3, 1]. For example, in applications where FM-based agents handle critical data, such as in medical diagnosis or or self-driving cars, incorrect outputs can lead to dangerous consequences [32]. Additionally, there is a risk of generating ethically or morally questionable content, which can damage the credibility and acceptance of the system [53].
@@ IV-A5 Fairness
FM-based agents can face bias and discrimination in model outputs. These biases can emerge from the training data, model algorithms, or deployment context [61, 2]. For instance, a language model used in recruitment for initial screening of CVs and responses to selection criteria might inadvertently favor candidates from certain demographics, cultures, and languages [26, 8], which ultimately undermines the credibility of the systems. Guardrails address these fairness issues and ensure equitable outcomes through model parameter adjustments. For instance, in our recruitment example, they can analyze the model’s decisions and correct potential biases. Furthermore, guardrails promote transparency and accountability by providing insights into how decisions are made [18, 1], and avoid harmful stereotypes [53].
@@ IV-A6 IP Protection
In FM-based agents, another key problem is the unauthorized use of generated content, where users might exploit outputs created by these systems without proper attribution or licensing, making them vulnerable to unauthorized use, duplication, or distribution [26, 55, 10]. Guardrails help mitigate these problems and better protect content’s copyright [49]. They also aid in compliance with copyright laws and licensing agreements to mitigate legal risks [19, 12]. Techniques such as watermarking, fingerprinting, and labeling ensure the copyright and ownership of data [10, 1].
IV-A7 Cost
FM-based agents can incur significant costs due to errors, inefficiencies, or non-compliance with regulations. Without proper guardrails, agents might generate outputs that lead to financial losses, legal penalties, or damage to reputation [26, 12]. For example, an agent that provides incorrect financial advice could result in monetary losses for users and potential lawsuits against the provider. Implementing guardrails helps mitigate these risks by preventing erroneous or harmful outputs, thereby reducing potential costs associated with rectifying mistakes, handling legal issues, or rebuilding trust [1, 62]. Additionally, guardrails can optimize resource utilization by preventing unnecessary computations or actions, leading to cost savings in system operations [53].
IV-A8 Performance
Performance is a critical concern in FM-based agents, as they are expected to provide timely and efficient responses [24]. Without appropriate guardrails, agents might engage in resource-intensive computations that degrade system performance or lead to unacceptable response times [54]. Guardrails help manage performance by monitoring and controlling the computational resources used by the agent [64, 53]. For instance, they can limit the complexity of tasks the agent undertakes or prevent it from entering endless loops, ensuring consistent and efficient operation [49]. Moreover, guardrails can enhance performance by filtering out irrelevant or low-quality inputs, allowing the agent to focus on processing meaningful data [16, 1].
IV-B Internal Quality Attributes for Multilayer Runtime Guardrails (RQ2)
We identified eight key quality attributes essential for the design of runtime guardrails in FM-based agents. Below, we discuss these attributes in detail, highlighting their roles and importance.
@@ IV-B3 Customizability
Customizable guardrails provide tailored protection that meets specific requirements and supports diverse operational needs in FM-based agents [1, 65]. They allow for adjustments and configurations that align with particular operational goals, data characteristics, and regulatory environments. For example, a customer service chatbot can enable priorities for different guardrails and adjust data handling based on the user’s location, ensuring compliance with GDPR in Europe and CCPA in the U.S. [66]. Customizability also includes the capability to integrate user-defined rules and policies, ensuring that the guardrails can support diverse operational contexts and evolving needs [64].
@@ IV-B6 Portability
Portability in guardrails for FM-based agents refers to the ability of these protective measures to be easily adapted and applied across different foundation model (FM) systems and platforms [27]. This includes ensuring that the guardrails function consistently across various FM architectures and environments, thereby maintaining their effectiveness and integrity regardless of the underlying system [26]. For example, the same guardrail can be applied for content moderation in both a customer service chatbot and a social media platform, regardless of their underlying technology. The benefits of designing portable guardrails include compatibility across multiple programming languages and frameworks facilitate their integration into diverse technological stacks [49]. These capabilities ensure that the guardrails remain effective and operational as the system evolves or migrates to new environments. Portable guardrails also support seamless updates and improve scalability to maintain high standards of security and compliance while adapting to new technological advancements within systems [16].
@@ IV-B8 Interpretability
IV-C Guardrails Design Options (RQ3)
This section presents a structured taxonomy for designing multi-layered runtime guardrails for FM-based agents, focusing on identifying various design alternatives.
@@ IV-C1 Guardrail Actions
Guardrail actions are crucial for addressing the specific needs of FM-based agents. We have identified the following guardrail actions as key elements for FM-based agents:
  • Block: The block action prevents specific inputs (such as user prompts) or outputs (such as content generated by FMs) from being processed or sent by various components (such as FMs and tools) in FM-based agents [54]. For example, the block action can reject the user prompts containing harmful instructions, thus preventing undesired outcomes.
  • Filter: The filter action involves scanning and removing undesired or irrelevant content from the inputs or outputs of different components in FM-based agents [69, 70]. For instance, a filter may remove any personal data contained in the user prompts or the output generated by FMs.
  • Flag: The flag action is used to mark specific inputs, outputs, or operations within FM-based agents [16]. For example, unusual transactions requested by the FM-based agent can be flagged for human review to ensure they comply with organizational policies [1, 30].
  • Modify: The modify action allows for the adjustment of inputs or outputs of various components in the FM-based agents to meet specific requirements or standards [9]. For example, the system can modify the user prompts by adding more context and examples, making it easier for the FM to accurately interpret the user’s intentions and provide more relevant responses.
  • Validate: The validate action checks inputs, outputs, and intermediate results against predefined criteria to ensure they meet specified requirements or standards [70, 26]. For example, the customer service chatbot could validate if the output generated by the FM-based system contains any sensitive company information.
@@
  • Simulate: The simulate action involves running tests in a controlled environment to predict and analyze potential outcomes [1]. For instance, a system might simulate different configurations to determine most effective plan.
@@
FM-based agents Prompts Block, Filter, Flag, Modify, Validate, Prioritize, Rate Limit, Parallel Calls, Retry, Defer, Isolate, Log
@@ IV-C3 Guardrails Scopes
The scope of guardrails in FM-based agents ranges from individual preferences to industry standards, reflecting the multi-layered approach of the Swiss Cheese Model. At the user level, guardrails reflect individual preferences and requirements. This involves adjusting the system’s behavior based on user-defined settings to align outputs with both user expectations and ethical considerations. Incorporating user preferences into guardrails provides a personalized experience while maintaining safety and compliance [55, 69]. Such guardrails ensure that the agents respects user autonomy and produces outputs that are relevant and acceptable.
At the organizational level, guardrails align with internal policies and procedures governing the operation and use of FM-based agents. This includes compliance with corporate governance, data protection policies, and ethical guidelines established by the organization [12]. Guardrails also ensure consistency and accountability across different departments and functions within the organization.
Industry-level regulations and standards provide the broader regulatory framework within which FM-based agents must operate. Guardrails designed to comply with these regulations guarantee that the system adheres to industry best practices and legal requirements [16]. Guardrails facilitate simpler auditing and certification processes, ensuring the system remains compliant with evolving regulatory landscapes.
@@ IV-C4 Rules
Guardrails rules can be configured in different ways: including uniform rules, priority-enabled rules, context-dependent rules, and negotiable rules, forming layers of defense as per the Swiss Cheese Model. A uniform strategy applies the same set of guardrails consistently across all scenarios, ensuring simplicity and uniformity [67]. It is particularly effective in environments with stable and well-understood risks. It largely reduces the complexity of managing diverse guardrails [55]. A priority-enabled strategy prioritizes certain guardrails based on the criticality and sensitivity of operations or data. Context-dependent strategies adjust the implementation of guardrails based on the system’s specific operational context. This allows for dynamic adjustments to guardrails in response to changing conditions, user needs, and operational environments [49]. The negotiability of guardrails, categorized into hard and soft, defines the level of flexibility in enforcing rules. Soft guardrails allow adjustments based on context and situational demands, providing a balance between protection and operational flexibility [49]. In contrast, hard guardrails are rigid and non-negotiable, ensuring adherence to critical legal, ethical, or safety standards [12, 32].
@@ IV-C6 Modality
The modality of guardrails refers to the types of data and interactions they manage. Guardrails can be designed for single modal or multimodal systems. Single modal systems operate with one type of data input or output, such as text, image, or audio. For instance, in text-based agent systems, guardrails focus on addressing issues like offensive language, misinformation, and data privacy [49]. In image-based agent systems, they may involve techniques for detecting explicit content or ensuring image quality standards [26].
@@ IV-C7 Underlying Technique
Rule-based models utilize predefined rules to monitor and control FM-based system behavior. These models implement strict and deterministic guidelines that the system must follow to ensure compliance with regulatory requirements for data access and processing [49]. They are particularly effective in environments where operational parameters are well-defined and stable. Rule-based agent systems can be updated and are somewhat flexible. However, they may still struggle with unexpected scenarios, such as detecting novel AI-generated content that falls outside predefined rules. This reliance on static rules can limit their adaptability and regular updates are needed [71, 16].
@@
Hybrid models integrate rule-based agent systems with the adaptability of machine learning models to respond to new threats and evolving data patterns [53]. For instance, Khorramrouz et al.[59] demonstrate the use of the PaLM 2 framework to process user input and dynamically implement rule-based decisions. This framework tests the system’s limits by iteratively generating toxic content to evaluate PaLM 2’s safety guardrails. However, integrating hybrid models can increase system complexity and create additional challenges [53].
V Discussion
In this section, we discuss the implications of our proposed taxonomy, its alignment with the Swiss Cheese Model for AI safety, and how it addresses the challenges in designing multi-layered runtime guardrails for FM-based agents. We also reflect on the potential impact of our findings on future research and practice in AI safety by design.
V-A Adapting the Swiss Cheese Model for Multi-Layered Guardrails
The Swiss Cheese Model, originally developed in risk management and safety science, illustrates how multiple layers of defense can prevent hazards from leading to adverse outcomes [75]. Each layer has potential weaknesses (holes), but when combined, they provide a robust barrier against failures. In the context of FM-based agents, the Swiss Cheese Model underscores the necessity of implementing multi-layered guardrails to address the unique challenges introduced by these agents’ autonomous and non-deterministic behaviors.
FM-based agents interact with various components during runtime, including goals, prompts, plans, tools, knowledge bases, and intermediate and final results. Each of these components represents a potential point of failure or risk. By adapting the Swiss Cheese Model, we propose that guardrails should be implemented at multiple layers of the agent architecture to effectively mitigate risks. This multi-layered approach ensures that if one layer fails to detect or prevent an undesirable behavior, subsequent layers can compensate, thereby enhancing the overall safety and reliability of the agent.
Implementing guardrails at different levels—such as input processing, model reasoning, action execution, and output generation—aligns with the Swiss Cheese Model’s principle of layered defenses, including overlapping layers. This approach addresses the complex interplay between different components in FM-based agents and provides a systematic framework for AI safety by design. It also highlights the importance of considering both external and internal quality attributes in guardrail design, as well as the trade-offs involved in different design options.
V-B Implications for Practice and Research
Our taxonomy offers practical guidance for developers and researchers in designing multi-layered runtime guardrails for FM-based agents. By categorizing the motivations, quality attributes, and design options, we provide a structured approach to address AI safety concerns effectively. Practitioners can use this taxonomy to make informed architectural decisions, balancing trade-offs between different quality attributes such as accuracy, performance, adaptability, interpretability and so on.
For researchers, our work assists in exploring new techniques and methodologies in guardrail design and implementation. The alignment with the Swiss Cheese Model encourages a holistic view of safety mechanisms, promoting research into how different layers can interact synergistically to enhance agent safety. Furthermore, understanding the relationships between guardrails, the risks they mitigate, and the quality attributes they influence can inform the development of standards and best practices in the field of AI safety.
V-C Threats to Validity
@@
Our study is subject to standard literature search and selection bias threats. We addressed these threats by searching the most commonly used databases in the IT and software engineering domains. We adjusted our search strings several times during the automatic search to maximize the number of relevant articles matching two key concepts: ‘guardrails’ and ‘FM-Based agents’. We also kept our search string generic to search through the titles, abstracts, keywords, and full text of articles to cover the maximum number of relevant papers. We then conducted a manual search on Google Scholar to complement the automatic search using a snowballing strategy. Furthermore, predefined review protocols with detailed inclusion and exclusion criteria helped us reduce bias in selecting primary studies. We applied several quality assessment criteria (shown in section III-F) to estimate the quality of the selected primary studies. Even though the proposed criteria were not too strict, applying them indeed caused several initially selected papers to be excluded. To mitigate the risk of missing important data from the primary studies, we reinstated the excluded papers closely related to the primary studies.
@@
Moreover, our definitions and categorizations may not capture all relevant aspects of guardrails in FM-based agents. To mitigate this threat, we validated the taxonomy through extensive literature review and expert feedback. However, this introduces a risk of producing biased results that address only expert needs, as the people involved in the feedback process have extensive experience in the AI and software engineering domains. Our review protocols helped us to reduce such bias.
@@
We prepared a guardrails taxonomy and conducted a comparative analysis of its components to help the reader better understand their design and evaluation. We critically examined the strength and consistency of relationships in the selected studies for a reliable taxonomy structure and drew conclusions. Nonetheless, the generalizability of guardrails to different contexts and types in FM-based agents remains a potential limitation, and specific adaptations might be necessary for certain systems, such as those used in healthcare or financial organizations. Additionally, our findings face a challenge as some of them may become outdated due to the rapid evolution of FM-based agents and associated guardrails.